OpenStudy (anonymous):

12x^(2)+2x-24 factor i need work shown step by step

OpenStudy (hoblos):

12x^2+2x-24 = 12x^2 + 18x -16x -24 = 6x(2x+3) -8(2x+3) = (2x+3)(6x-8)